OpenStudy (anonymous):

what's the derivative of [(1-(x/9)]????⊙_⊙

OpenStudy (anonymous):

=_=im stuck here..

OpenStudy (amistre64):

stuck on which part?

OpenStudy (anonymous):

x/9...

OpenStudy (amistre64):

if i asked you what the derivative of 5x was, what would you conclude?

OpenStudy (amistre64):

im assuming you know the power rule ....

OpenStudy (anonymous):

Mmmm..i forgot..1/5x????????

OpenStudy (anonymous):

ohhhhhhhhh

Parth (parthkohli):

Well, \(\dfrac{x}{9}\) can be written as \(cx\). What is \(c\)?

OpenStudy (amistre64):

\[\frac{d}{dx}kx^n=kn~x^{n-1}\] \[\frac{d}{dx}5x^1=5(1)~x^0=5\]

OpenStudy (amistre64):

\[\frac{d}{dx}\frac19x^1=?\]

Parth (parthkohli):

Oh, \(c\) has been disclosed by Amistre. :-)

OpenStudy (amistre64):

<<<raises hand and waves it vigorously .... pick me pick me, i know what c is!!!

OpenStudy (anonymous):

c is 1/9;D

Parth (parthkohli):

Right there.

Parth (parthkohli):

I can do this without the power rule, just basic algebra. :-P

OpenStudy (amistre64):

so, what is the deriative of say:\[y=x^{2x}\]

Parth (parthkohli):

\[y = -\dfrac{1}{9}x + 1 \]As they say, the derivative is just the slope. What is the slope of this line I gave ya above?

Parth (parthkohli):

No, that one is Chain Rule. :-)

OpenStudy (anonymous):

@-@IM sorrrrrry my internets working 2 bad!!!!ill @ u after 10minn

OpenStudy (anonymous):

Is it 1/9x?????????????????????????@p

OpenStudy (anonymous):

sorry @ParthKohli

Parth (parthkohli):

Hmm, the slope of \(y = mx + b\) is only \(m\)

OpenStudy (anonymous):

⊙△⊙?

OpenStudy (anonymous):

oooohhhh goshhhhhhhhh finally got it!!!thanks soooo much~!>A<@ParthKohli

Parth (parthkohli):

:-)